OpenStudy (geometry_hater):

Solve for X using quadratic function 3y^2 - 2y = 5 I already subtracted the 5 to get 0 so my A = 3^2 b= -2 c= -5 But im having trouble with the actual equation part which happens to be -2 + or - square root -2^2 - 4 * 9 * -5 over 2 * 9.

OpenStudy (anonymous):

\[3y ^{2}-2y-5\] \[-(-2)\pm \sqrt{(-2)^{2}-4(-3)(-5)}/(2(3))\] your mistake is squaring the first coefficient. You do not need to. Recalculate from here, it should work out now.

OpenStudy (anonymous):

sorry also -(-2) is positive 2 as well.

OpenStudy (geometry_hater):

ok so my A coefficient is just 3

OpenStudy (anonymous):

Correct

OpenStudy (geometry_hater):

Ok now what do i do in the equation?

OpenStudy (anonymous):

I also made the mistake of -3 in the equation it is actually 3

OpenStudy (anonymous):

I'll equate it out from the original

OpenStudy (anonymous):

\[(2\pm \sqrt{64})/6\] \[2\pm8/6\] =2-8=-6/6=-1 or 10/6=5/3

OpenStudy (anonymous):

y=

OpenStudy (geometry_hater):

Thanks